OpenStudy (anonymous):

Are each of the statements below true or false? Explain your answer without just saying true or false. o √a + √b = √(a + b) Explain why. o The numerator and denominator of the following must be multiplied by √3 to rationalize 3/(3 + √3). Explain why.

myininaya (myininaya):

\[\sqrt{a}+\sqrt{b} \neq \sqrt{a+b}\]

OpenStudy (moongazer):

false, because √(a + b) means that you should add first a and b before having it √. unlike in √a + √b you will only add the √ of a and b

myininaya (myininaya):

counterexample: if a=9 and b=25, then \[\sqrt{9}+\sqrt{25}=3+5=8 \neq \sqrt{34}=\sqrt{9+25}\]

myininaya (myininaya):

second one: you must multiply the fraction by

myininaya (myininaya):

\[\frac{3-\sqrt3}{3-\sqrt3}\]

OpenStudy (anonymous):

myininaya so the second one is false too right

myininaya (myininaya):

yes it is false you change the value of that fraction when you multiply by square root of 3 you need to multiply by \[1=\frac{3-\sqrt{3}}{3-\sqrt{3}}\]

OpenStudy (anonymous):

myininaya thank you very much

myininaya (myininaya):

np